OpenStudy (anonymous):

Write an equation that passes through (0, -2) and whose slope is ½. y = 1/2x + 2 y = -1/2x – 2 y = 1/2x – 2 y = -2x + 1/2 ?????

OpenStudy (callisto):

This may help you: \[y-y_1 = m(x-x_1)\], where \((x_1, y_1) \) are the coordinates of a point lying on the line and m is the slope of the line.

OpenStudy (phi):

If you have an equation in the form y=mx+b then you have the equation of a line in slope-intercept form m (the coefficient of x) is the slope of the line b is the y-intercept (the y value when x is 0) notice that your problem gives the slope of the line and the y-intercept ( the 2nd number in the (x,y) pair (0,-2) )

OpenStudy (anonymous):

It is either one or three. for x=0, y=-2 so the answer is y = 1/2x – 2